Assume inflation in Turkey is 50% and 2% in the US. Using the relative PPP, 


 a) Calculate the exchange rate (R= TL/$) and the exchange rate (R= $/TL).  

b) Assume that the Turkish inflation  increases to 100% while the US  inflation remains at 2%, calculate the  new exchange rate between the Turkish Lira and the US dollar.


1
Expert's answer
2022-04-11T12:09:57-0400

1. Changes in exchange rate:

"e_f=\frac{1+i_h}{1+i_f}-1"

"e_{TL}=\frac{1+0.02}{1+0.5}-1=-0.32"

"e_{USD}=\frac{1+0.5}{1+0.02}-1=0.47"

2. Current exchange rate is 14.75 TL/USD. So, the new exchange rate will be:

"S_t=S_0\frac{1+i_A}{1+i_B}-1=14.75*\frac{1+1}{1+0.02}=28.92"